Launch a tennis ball at 120 mph, from the ground level. You get horizontal range of 115 m. Find

the initial launch angle.

Launch a tennis ball at 100 mph, from the ground level. You get a maximum height of

50 m. Find

the initial launch angle.

explain steps!!!
Answer #1

A)

120 mph = 53.64 m/s

Range, R = u^2 sin(2 theta)/g

115 = 53.64^2 sin(2 theta)/9.8

Theta = 11.53°

B)

100 mph = 44.7 m/s

H = (u sin(theta))^2/2g

50 = (44.7 sin(theta))^2/(2 x 9.8)

(44.7 sin(theta))^2 = 50 x 2 x 9.8 = 980

Taking square root both sides,

44.7 sin(theta) = 31.30

Sin(theta) = 0.7

Theta = arcsin(0.7)

Theta = 44.45°

Comment in case any doubt please rate my answer....